OpenStudy (anonymous):

A= (f+gDiv by 2)h solve for g how can i solve for g?

OpenStudy (aum):

Which one?\[ A = (f + \frac g2)h ~~~~\text{ OR }~~~~ A = \left(\frac{f+g}{2}\right)h \]

OpenStudy (anonymous):

second one

OpenStudy (aum):

\[A = \left(\frac{f+g}{2}\right)h \\ \text{Multiply by 2:} \\ 2A = (f+g)h \\ \text{Divide by h:} \\ \frac{2A}{h} = f + g \\ \text{Subtract f:} \\ \frac{2A}{h} - f = g \\ g = \frac{2A}{h} - f = \frac{2A-fh}{h} \]

OpenStudy (anonymous):

is that it? Or is there more?